# Dual-Agent Reinforcement Learning for Adaptive and Cost-Aware Visual–Inertial Odometry

### v1

@@ D. Additional Experimental Results
Additional baseline: Visual-Selective-VIO on EuRoC.
To further assess adaptive VIO methods trained on driving data, we evaluate Visual-Selective-VIO (VS-VIO) on EuRoC without any retraining. We reuse the official KITTI-pretrained model and convert EuRoC sequences MH_02–MH_05 into the KITTI-style format used by the original code, including nearest-neighbor timestamp alignment between images and ground-truth and IMU interpolation as described in Sec. 4.1. Table 10 reports the SE(3)-aligned RMSE ATE (m). Compared to both classical VIO and our method (see Tables 1 and 8), VS-VIO exhibits substantially larger trajectory errors on these aggressive 6-DoF MAV sequences, confirming a strong domain gap between ground-vehicle training data and EuRoC. We therefore treat VS-VIO as a conceptually related method, but do not include it as a main numerical baseline in the core results tables.
E. Limitations and Future Work

### v2

@@ A.2. PPO Hyperparameter Summary
The training stability of our dual-agent system is ensured through specific strategies tailored to each agent’s reward structure:
  • Select Agent: Due to sparse rewards and high penalties for incorrect skipping, the training process is naturally sensitive. We stabilized convergence by employing a higher entropy coefficient to encourage exploration and a conservative learning rate, preventing premature convergence to sub-optimal policies.
  • Fusion Agent: We utilize a Curriculum Learning approach. Before RL training, MLP1 undergoes supervised pre-training. This provides robust velocity priors, avoiding the instability often associated with end-to-end reinforcement learning from scratch.
@@ D. Additional Experimental Results
Additional baseline: Visual-Selective-VIO on EuRoC. To further assess adaptive VIO methods trained on driving data, we evaluate Visual-Selective-VIO (VS-VIO) on EuRoC without any retraining. We reuse the official KITTI-pretrained model and convert EuRoC sequences MH_02–MH_05 into the KITTI-style format used by the original code, including nearest-neighbor timestamp alignment between images and ground-truth and IMU interpolation as described in Sec. 4.1. Table 10 reports the SE(3)-aligned RMSE ATE (m). Compared to both classical VIO and our method (see Tables 1 and 8), VS-VIO exhibits substantially larger trajectory errors on these aggressive 6-DoF MAV sequences, confirming a strong domain gap between ground-vehicle training data and EuRoC. We therefore treat VS-VIO as a conceptually related method, but do not include it as a main numerical baseline in the core results tables.
Evaluation on failure modes. We conduct stress tests on the EuRoC dataset (MH_04) under three extreme conditions: (1) VO Outage, (2) Severe Blur, and (3) IMU Noise. The results in Table 11 demonstrate that while accuracy degrades under extreme stress, our method significantly outperforms the DPVO baseline. This resilience is primarily attributed to the Fusion Agent, which dynamically detects low visual reliability and suppresses visual weights, effectively leveraging the IMU Bias Estimator as a safety net.
Despite the improved robustness, residual drift in VO-denied areas suggests two directions for future work: (1) incorporating long-range temporal context into confidence estimation, and (2) establishing a feedback mechanism to re-initialize the VO backend using IMU states during prolonged outages.
Table 11: Robustness Stress Test under Severe Sensor Degradation.
Method Nominal VO Outage Severe Blur IMU Noise
(Clean Data) (Blackout 2s) (30% Frames) ($2\times$ Noise)
DPVO 0.106 1.928 1.122 -
Ours 0.092 1.114 0.472 0.128
Scale Robustness and Monitoring. Regarding scale estimation, we acknowledge that our base MLP1 primarily performs velocity smoothing and has limited authority to correct large initial scale errors. To address this, we integrated a lightweight Online Scale Monitor. This module calculates the motion magnitude ratio between IMU pre-integration and VO over a sliding window (15 frames). If a persistent deviation is detected, a correction factor realigns the VO scale with the IMU’s metric reference. As shown in Table 12, under $\pm 20\%$ initialization error, our system with the Monitor effectively mitigates the scale drift.
Table 12: Scale Robustness Test (ATE [m]).
Method Nominal Small Init Large Init
Ours (Base) 0.092 0.661 0.682
Ours (+Monitor) 0.096 0.522 0.508
Generalization: To assess performance with minimal domain adaptation, we evaluated a model fine-tuned only on TUM-VI (Corridor_4) applied to EuRoC (MH_04). ATE increases (0.092m $\to$ 0.173m) due to the domain gap in IMU noise profiles. However, the system successfully completes the trajectory, validating robustness to unseen sensors.
E. Discussion on Algorithmic Novelty
To further clarify our contributions relative to existing RL-based Visual Odometry (VO) works, we emphasize two fundamental distinctions in our dual-agent architecture:
1. Paradigm Shift: “Shift Left” Pre-emptive Gating. Traditional RL-based VO methods typically focus on post-processing, such as filtering keyframes or weighting features after they have been extracted. This represents a sunk cost where significant computation has already been spent. In contrast, our Select Agent proposes a “Shift Left” strategy: it performs pre-emptive gating on raw sensor data streams. By making skipping decisions before the visual frontend is invoked, we bypass the entire tracking and bundle adjustment pipeline for redundant frames.
2. First RL-based VIO Fusion. While prior works often use RL for visual parameter tuning, our framework is, to the best of our knowledge, the first to employ RL for dynamic IMU-Visual fusion. Our Fusion Agent goes beyond simple weighting; it learns a context-dependent policy that adapts the IMU-visual contribution based on motion dynamics and visual reliability.
F. Limitations and Future Work